# LaserEscape: Detecting and Mitigating Optical Probing Attacks

### v1

@@ 5.1. Moving Target Defense via PR
Figure 6 illustrates the design of the proposed MTD-based response for the optical attack on an FPGA. We consider an original hardware design which includes some sensitive data requiring protection. For simplicity, we assume the key registers $R_{1}$, $R_{2}$, and $R_{3}$ as the secret elements in our design. Using a the TCL language, the targeted modules are specified by the user through predefined annotations in register-transfer level (RTL). At this stage, the static design in the original bitstream as well as sample partial files are generated. The constraints such as the allowable range for FPGA slices, regional locations, and possible range of flip-flops (FFs) (Xilinx, 2023b) are established for generating reconfiguration. This information, together with the generated bitstreams, are moved to the online phase of the framework. In this phase, a lightweight processing unit (PS) is employed within the SoC (or in an external processor) to execute the reconfiguration. A secure, one-time pseudo random number generator (PRNG) (Tsoi et al., 2003; Monfared et al., 2020) is implemented to provide randomness to the PR-generator unit each time. The PR-generator includes a bitstream manipulator (Byteman(Manev et al., 2022) in our instance), uses the predefined constraint limits to randomized location (LOC) and shuffle the chosen registers. These two randomization introduce coarse and fine-grain randomness to the reconfiguration, respectively. Consequently, the corresponding partial reconfiguration bitstream is created based on the original bitstream. The FPGA is then reprogrammed as soon as the trigger signal from the sensor is received.
@@ 5.2. Gate Polymorphism
In general, for a n-input LUTs, we can implement a polymorphic n-1-input LUTs using a controller input to change the logic function. For instance, we can exploit input $i3$ in Fig. 7(b) to switch the output between 1) The result of a 3-input AND and 2) constant zero. This idea can be generalized to implement other logic gates (i.e., XOR) easily by selecting proposer LUT inputs. Furthermore, by connecting the control input pin to the latch output of the sensor described in Sect. 4.2, we can modify any protected logic upon detecting optical probing attacks.
@@ 6. Experimental Setup
Electrical and Optical Equipment. A Hamamatsu PHEMOS-X FA microscope was used to perform optical probing, see Fig. 8. It is equipped with a 1.3 $\mu$m high-power laser source used for our analysis. Additionally objective lenses of 5X/0.14NA, 20X/0.6NA, 50X/0.76NA, and 71X are employed in our experiments. Furthermore, optical zooms of 2x, 4x and 8x were higher magnifications. As illustrated in Figure 8(a), in the EOFM process, the DUT is scanned by a laser through the use of galvanometric mirrors.
@@
Figure 9. EOFM probing attack timeline.
@@
Communication. For the run-time configuration/communication of the target FPGA, we used a 3.3V modified Arduino Uno R3 communicating over I2C interface with the FPGA target. A compact circuitry on the FPGA part is deployed to handle I2C communication and commands from the user. We choose an I2C bus since the addressable communicated data makes adding circuits easy for automated experiments.
@@
Analyzer and Controller Configuration. Illustrated in Figure 8(a), we use a dedicated controller system, to send commands, program key registers, adjusting target clock, and handle I/O for experiments via I2C. The measurement system is a dedicated system synced with the PHEMOS-X to control the stage, laser configurations and display the digitized data captured from the detection RAC. We use Semishop Failure Analysis Software 4.0 to control the stage and capture the images and signals.
@@
Sensor Configuration. After implementing the target RTL, we added the sensor in a neighbour SLICE of the target. In our sensor implementation, we used a chain of 8-long IDELAYE2 ($m=8$) for clock delay path. The tuning time interval is set to $t_{sense}=100ms$ and the zero counter interval window in test time is set to $t_{detect}=255$ clock cycles.
@@ 7.1. EOFM for Localizing Target Registers
7.2. EOP for Registers Values
Upon successful localization of the registers, we can then mount an EOP attack by focusing the HIL beam on each target FF. This process can be executed during the secret operation of the IC. For instance, if the state registers of a block cipher (e.g., AES) are probed, the keys are revealed easily. Fig 10 showcases a successful EOP attack on an implementation of shift-register.
@@ 7.3. EOFM for Function Recovery
As an extension to the key recovery attempt, we also demonstrate a function recovery attack via EOFM. Similar to previous probings, the adversary localizes the region of interest by exploiting the control signals in the frequency domain. By conducting a careful profiling procedure, a reverse-engineering approach is possible to discover the functionality of the building blocks. Fig. 11 serves as an example for a series of profiling EOFM patterns of a function. It can be inferred that the correlation of the FFs patterns can be described by an XOR function among the 8 input FFs at the left side of the slice and the output FFs at the right of the target slice.
@@ 7.4. Evaluating Sensor Trigger
In order to unsure the sensor operates properly, we empirically investigate the output of the sensor. As an example, Fig 12 displays the normalized zero counter output of the sensor in two test cases during a time window of $t_{detect}=255$ clock cycles. As it can be observed, while the laser is activated, the average of the zero counter samples is significantly larger compared to the scenario where the laser is idle. Furthermore, Fig 12 shows that in this case, the trigger threshold can be selected by a safe margin.
@@ 7.5. Defeating Function Discovery Attacks
For this experiments, we deploy LaserEscape sensor on the target FPGA. We mount the same optical probing attack described in Section 7.3. In this experiments instead of a simple XOR gates for the target function we deployed the proposed polymorphic XOR gates in the target design. For the EOFM, Figure 13 displays the input and output registers of the target function. As shown in the upper part of Figure 13, the unprotected target reveals the XOR function for $in_{1}=0b0101$ and $in_{2}=0b1010$ which results in $out=0b1111$ highlighted on the right side of the figure. When sensor is activated on the FPGA, once the laser emission is detected, the polymorphic gate is activated and cleared the output of the XOR gates. This is highlighted in the lower part of the Figure 13.
@@ 7.6. Defeating Register Recovery
To evaluate our method we organize the similar EOFM attack described in Sect. 7.1 in presence of LaserEscape. We target a set of 8 FDRE registers which are considered to be protected in the design. One can assume these registers hold the master key of a cryptographic algorithm (e.g., AES). Fig. 9(b) illustrated the timeline of the EOFM scan targeting the FDRE registers. Compared to the similar attack depicted in Fig.9(a), as the scanning laser line approaches the physical locations of the target FDREs, the sensor in LaserEscape is triggered and consequently the MTD via PR is carried out in real-time, leaving no trace of the FDRE targets once the laser arrives at the target destination.
@@ 8. Discussions
Overhead Analysis. In terms of the overhead introduced by the LaserEscape, the 1LUTsensor is very efficient in terms of area overhead by employing a single LUT and utilizing 9 IDELAYE2 in the design, which is negligible. For the MTD, if 32-bit state registers of an AES block cipher is considered as the target, a total area overhead of 3.5% (compared to original AES) is estimated by Vivado toolkit in the case of choosing 8 Reconfigurable LOCs. The delay overhead caused by the PR process for the same implementation is measured to be $223\mu s$, which maintains a large safe gap against the laser dwell time per pixel for EOP and EOFM attacks.
Stability. As long as the ambient conditions and FPGA load remain the same, our sensor remains stable. We let the sensor run idle for about 21 hours, logging the sensor’s readings over zero counts. Our experiments show that the maximum zero count had plateaued after less than 20 minutes, but the rolling average stayed at around zero for the rest of 21.5 hours. This implies that the zero counts follow a normal distribution and the sensor is stable.
Further Improvements It is possible to statistically investigate the area coverage of the delay sensor with regards to emitted laser power, scan duration, corresponding connections of the sensor circuitry (e.i, wires), and different tuning of the IDELAYs or any similar delay-based sensor against optical attacks. It would also be worthy to study the possible improvement of detection area coverage if multiple sensors are deployed in the target circuit.
@@ Acknowledgment
This effort was sponsored by NSF under the grant numbers 2150123 and 2150122.

### v2

@@ 5.1. Moving Target Defense via PR
Fig. 6 illustrates the design of our proposed MTD-based response for an optical attack on an FPGA. We consider an original hardware design which includes some sensitive data requiring protection. For simplicity, we assume the key registers $R_{1}$, $R_{2}$, and $R_{3}$ as the secret elements in our design. Using TCL, the targeted modules are specified by the user through predefined annotations in register-transfer level (RTL). At this stage, the static design in the original bitstream as well as sample partial files are generated. The constraints such as the allowable range for FPGA slices, regional locations, and possible range of flip-flops (FFs) (Xilinx, 2023b) are established for generating reconfiguration. This information, together with the generated bitstreams, are moved to the online phase of the framework. In this phase, a lightweight processing unit (PS) is employed within the SoC (or in an external processor) to execute reconfiguration. A secure, one-time pseudo random number generator (PRNG) (Tsoi et al., 2003; Monfared et al., 2020) is implemented to provide randomness to the PR-generator unit each time. The PR-generator includes a bitstream manipulator (Byteman(Manev et al., 2022) in our instance), uses the predefined constraint limits to randomized location (LOC) and shuffle the chosen registers. These two randomization introduce coarse and fine-grain randomness to the reconfiguration, respectively. Consequently, the corresponding partial reconfiguration bitstream is created based on the original bitstream. The FPGA is then reprogrammed as soon as the trigger signal from the sensor is received.
@@ 5.2. Gate Polymorphism
In general, for $n$-input LUTs, we can implement polymorphic $n-1$-input LUTs using a control input to change the logic function. For instance, we can exploit input $i3$ in Fig. 7(b) to switch the output between 1) The result of a 3-input AND and 2) constant zero. This idea can be generalized to implement other logic gates (i.e., XOR) easily by selecting proposer LUT inputs. Furthermore, by connecting the control input pin to the latch output of the sensor described in Sect. 4.2, we can modify any protected logic upon detecting optical probing attacks.
@@ 6. Experimental Setup
Electrical and Optical Equipment. A Hamamatsu PHEMOS-X FA microscope was used to perform optical probing, see Fig. 8. It is equipped with a 1.3 $\mu$m high-power laser source used for our analysis. Additionally objective lenses of 5X/0.14NA, 20X/0.6NA, 50X/0.76NA, and 71X are employed in our experiments. Furthermore, optical zooms of 2x, 4x and 8x were higher magnifications. As illustrated in Fig. 8(a), in the EOFM process, the DUT is scanned by a laser through the use of galvanometric mirrors.
@@
Figure 9. EOFM probing attack timeline on images captured during laser scanning from back-side of the FPGA.
@@
Communication. For the run-time configuration/communication of the target FPGA, we used a 3.3V modified Arduino Uno R3 communicating over I2C interface with the FPGA target. A compact circuitry on the FPGA part is deployed to handle I2C communication and commands from the user. We choose an I2C bus since the addressable communicated data makes adding post-processing circuits easy for different experiments.
@@
Analyzer and Controller Configuration. Illustrated in Fig. 8(a), we use a dedicated controller system, to send commands, program key registers, adjusting target clock, and handle I/O for experiments via I2C. The measurement system is a dedicated system synced with the PHEMOS-X to control the stage, laser configurations and display the digitized data captured from the detection RAC. We use Semishop Failure Analysis Software 4.0 to control the stage and capture the images and signals.
@@
Sensor Configuration. After implementing the target RTL, we added the sensor in a neighbour SLICE of the target. In our sensor implementation, we used a chain of 8-long IDELAYE2 ($m=8$) for clock delay path. The tuning time interval is set to $t_{sense}=100ms$ and the zero counter interval window in test time is set to $t_{detect}=255$ clock cycles. We chose $m$ and $t_{detect}$ because they were the largest values we could read or write with a single byte over I2C bus. We chose $t_{sense}$ because to achieve a balance of sensitivity versus speed when conducting experiments.
@@ 7.1. EOFM for Localizing Target Registers
Images captured from back-side FPGA in Fig. 9(a) show how the scanning is carried out (from left to right). As the horizontal scan line (across Y axis) moves, details of the internals of the FPGA are shown. The bright spots represent the electronic elements (e.g., registers) that actively operate on particular frequencies within the range of the EOFM scan. The target register areas, highlighted in a red box, are programmed registers that are localized by profiling. Similar to methods in (Lohrke et al., 2016), the profiling process consists of 1) programming the elements to fluctuate in a certain frequency, 2) activate/deactivate them and 3) analyze the optical activity extracted from EOFM. In our case shown in Fig. 9(a), 1) the target FDRE registers are programmed to reset with EOFM scanning frequency, 2) are set and clear in multiple iterations and 3) are inspected visually after EOFM scan in terms of the optical activity. The bright area within the red target area indicates that 1) the registers are fluctuating on the particular frequency and 2) they are set to one.
7.2. EOP for Register Values
Upon successful localization of the registers, we can then mount an EOP attack by focusing the HIL beam on each target FF. This process can be executed during the secret operation of the IC. For instance, if the state registers of a block cipher (e.g., AES) are probed, the keys are revealed easily. Fig 10 showcases a successful EOP attack on an implementation of shift-register. In this particular scenario, we focus the lens onto the registers (which are already localized via EOFM) and carry out the shift operation that is programmed on the FPGA. The probing process is synchronized with the operations on the FPGA via triggers. The probing experiment is performed during a particular interval with picosecond time accuracy to capture fast logical operations which results in voltage alterations in nodes on the chip. After multiple iterations and measurement averaging, the curves illustrated in Fig 10 are produced. Sometimes signal processing techniques, such as frequency band filtering and de-noising, are required to extract the clean states of the targets. This is due to the errors in placing and focusing the lens at the target element. As depicted in Fig 10, the shift register operations in REG 1 and REG 2 can be followed synchronously with the CLK-BUF rising edges.
@@ 7.3. EOFM for Function Recovery
As an extension to the key recovery attempt, we also demonstrate a function recovery attack via EOFM. Similar to previous probings, the adversary localizes the region of interest by exploiting the control signals in the frequency domain. By conducting a careful profiling procedure, a reverse-engineering approach is possible to discover the functionality of the building blocks. Fig. 11 serves as an example for a series of profiling EOFM patterns of registers in a function. It can be inferred that the correlation of the FFs patterns can be described by an XOR function among the 8 input FFs at the left side of the image and the output FFs at the right of the image. As indicated, the programmed functionality for this function is a simple bit-wise XOR operation where $c_{i}=a_{i}\oplus b_{i}$. By analysing the EOFM optical response from different input patterns, we can visually confirm the activation patterns for each individual register bits. By a simple observation at the first two rows of Fig. 11, one can suggest that the response of the neighbouring underlying registers are separated two by two. Particularly as highlighted by red squares, the optical patterns of $a_{3}a_{2}$ and $a_{1}a_{0}$ are physically separated. This might suggest the physical realizations of FFs in each slice on this particular FPGA architecture. Furthermore by analyzing the last row in Fig. 11, we can recover the functionality of the operations programmed for this target. Specifically, for $a_{1}a_{0}=0b01$, we can observe that that brightness pattern differs from $110b$ 11. With the same observation on $c_{1}c_{0}=0b10$ we can recover the functionality of the underlying operation in this analysis.
@@ 7.4. Evaluating Sensor Trigger
In order to ensure the sensor operates properly, we empirically investigate the output of the sensor. As an example, Fig 12 displays the normalized zero counter output of the sensor in two test cases during a time window of $t_{detect}=255$ clock cycles. As it can be observed, while the laser is activated, the average of the zero counter samples is significantly larger compared to the scenario where the laser is idle. Furthermore, Fig 12 shows that in this case, the trigger threshold can be selected by a safe margin.
@@ 7.5. Defeating Function Discovery Attacks
For this experiment, we deploy LaserEscape sensor on the target FPGA. We mount the same optical probing attack described in Sect. 7.3. In this experiment instead of a simple XOR gates for the target function we deployed the proposed polymorphic XOR gates in the target design. For the EOFM, Fig. 13 displays the input and output registers of the target function. As shown in the upper part of Fig. 13, the unprotected target reveals the XOR function for $in_{1}=0b0101$ and $in_{2}=0b1010$ which results in $out=0b1111$ highlighted on the right side of the figure. When the sensor detects laser emission on the FPGA, the polymorphic gate activates and clears the output of the XOR gates. This is highlighted in the lower part of Fig. 13.
@@ 7.6. Defeating Register Recovery
To evaluate our method we organize the similar EOFM attack described in Sect. 7.1 in presence of LaserEscape. We target a set of 8 FDRE registers which are considered to be protected in the design. One can assume these registers hold the master key of a cryptographic algorithm (e.g., AES). Fig. 9(b) illustrated the timeline of the EOFM scan targeting the FDRE registers. Compared to the similar attack depicted in Fig.9(a), as the scanning laser line approaches the physical locations of the target FDREs, the sensor in LaserEscape is triggered and consequently the MTD via PR is carried out in real-time, leaving no trace of the FDRE targets once the laser arrives at the target destination. Note that the sensor circuitry is placed on a neighbouring slice to the target shown in Fig. 9(b) and is not illustrated in these captures.
@@ 8. Discussions
8.1. Overhead Analysis
In terms of the overhead introduced by LaserEscape, the 1LUTsensor is very efficient in terms of area overhead by employing a single LUT and utilizing 9 IDELAYE2 in the design, which is negligible. For the MTD, if 32-bit state registers of an AES block cipher is considered as the target, a total area overhead of 3.5% (compared to original AES) is estimated by Vivado toolkit in the case of choosing 8 Reconfigurable LOCs. The delay overhead caused by the PR process for the same implementation is measured to be $223\mu s$, which maintains a large safe gap against the laser dwell time per pixel for EOP and EOFM attacks.
8.2. Stability
As long as the ambient conditions and FPGA load remain the same, our sensor remains stable. We let the sensor run idle for about 21 hours, logging the sensor’s readings over zero counts. Our experiments, shown in Fig. 14, show that the maximum zero count had plateaued after less than 20 minutes, but the rolling average stayed at around zero for the rest of 21.5 hours. This implies that the zero counts follow a normal distribution and the sensor is stable.
Figure 14. Plot of maximum and rolling average zero counts over 21.5 hours
8.3. Further Improvements
It is possible to statistically investigate the area coverage of the delay sensor with regards to emitted laser power, scan duration, corresponding connections of the sensor circuitry (e.i, wires), and different tuning of the IDELAYs or any similar delay-based sensor against optical attacks. It would also be worthy to study the possible improvement of detection area coverage if multiple sensors are deployed in the target circuit.
@@ Acknowledgment
This effort was sponsored by NSF Grants CNS-2150123 and CNS-2150122.